# A Cartesian Cut-Cell Two-Fluid Method for Two-Phase Diffusion Problems

### v1

@@ Abstract
We present a Cartesian cut-cell finite-volume method for sharp-interface two-phase diffusion problems in static geometries. The formulation follows a two-fluid approach: independent diffusion equations are discretized in each phase on a fixed staggered Cartesian grid, while the phases are coupled through embedded interface conditions enforcing continuity of normal flux and a general jump law. Cut cells are treated by integrating the governing equations over phase-restricted control volumes and faces, yielding discrete divergence and gradient operators that are locally conservative within each phase. Interface coupling is achieved by introducing a small set of interfacial unknowns per cut cell on the embedded boundary; the resulting algebraic system involves only bulk and interfacial averages. A key feature of the method is the use of a reduced set of geometric information based solely on low-order moments (trimmed volumes, apertures and interface measures/centroids), allowing robust implementation without constructing explicitly cut-cell polytopes. The method supports steady (Poisson) and unsteady (diffusion) regimes and incorporates Dirichlet, Neumann, Robin boundary conditions and general jumps. We validate the scheme on one-, two- and three-dimensional mono- and diphasic benchmarks, including curved embedded boundaries, Robin conditions and strong property/jump contrasts. The results demonstrate the expected convergence behavior, sharp enforcement of interfacial laws and excellent conservation properties. Extensions to moving interfaces and Stefan-type free-boundary problems are natural perspectives of this framework.
@@ 1 Introduction
Diffusive transport across material interfaces is central to a wide range of heat- and mass-transfer phenomena, including conjugate heat transfer between solids and fluids, interphase species transfer, dissolution/precipitation in porous media and phase-change processes. At the continuum level, these problems are naturally formulated as two-phase diffusion models in which a scalar field (e.g. temperature or concentration) satisfies diffusion equations with phase-dependent coefficients and is coupled across a sharp interface through jump relations and flux balance (bird_transport_2007; cussler_diffusion_2009; carslaw_conduction_1980; Ishii2011). Depending on the application, the interface may enforce continuity of the scalar (conjugate heat transfer), a weighted jump such as Henry’s law (partitioning in mass transfer) or a finite interfacial resistance such as Kapitza resistance (Sander2023; pollack_kapitza_1969). These interfacial relations must be satisfied accurately to predict transfer rates, which are often controlled by gradients localized near the interface.
From a numerical standpoint, sharp-interface diffusion on complex geometries is challenging whenever the computational mesh does not conform to the interface. Body-fitted (ALE-type) approaches can represent the interface explicitly but require mesh generation and mesh motion, with the attendant complexity and robustness issues in three dimensions (hirt_arbitrary_1974). On fixed Cartesian grids, a broad class of non-conforming strategies has therefore been developed. Immersed-boundary and ghost/immersed-interface-type methods modify stencils near the interface or introduce forcing terms so that boundary and interface conditions are satisfied approximately, often with excellent accuracy for smooth solutions but with less direct control over strict local conservation (peskin_flow_1972; gibou_second-order-accurate_2002; gabbard_high-order_2024). Interface-capturing methods (VOF or level-set approaches) (hirt_volume_1981; osher_fronts_1988; popinet_accurate_2009) coupled with a one-fluid formulation offer flexibility for complex topological changes; however, enforcing two-phase diffusion jump conditions sharply on a fixed grid typically requires additional reconstruction and coupling machinery (and, in practice, often relies on one-fluid effective-property closures in interfacial cells). Hybrid combinations have recently been explored for solidification/melt problems (limare_hybrid_2023) and XFEM/level-set formulations have also been used for coupled Stefan/mass-transport settings (li_numerical_2019), highlighting continued interest in sharp and accurate interfacial transfer.
@@
  • a validation on one-, two- and three-dimensional benchmarks, mono- and diphasic, demonstrating the expected accuracy behavior together with conservation properties at the discrete level.
The paper is organized as follows. Section 2 introduces the two-phase diffusion model and the interfacial jump conditions considered. Section 3 presents the embedded-boundary geometric setting and the reduced moment description used throughout. Section 4 derives the cut-cell finite-volume operators and the coupled discrete system. Numerical results are then reported in Section 5 to assess accuracy, robustness and conservation and Section 6 concludes and outlines extensions.
@@ 2 Continuum modeling
We introduce a generic scalar field $\phi$ (e.g. temperature $T$ or concentration $c$). The following describes the bulk transport model and the interfacial matching conditions used to close the two-phase diffusion problem.
@@ 2.1 Bulk transport equations
where $K^{\pm}$ is the diffusive mobility. The capacity $C^{\pm}$ collects the storage term (e.g., $C^{\pm}=\rho^{\pm}c_{p}^{\pm}$ and $K^{\pm}=k^{\pm}$ for heat; $C^{\pm}=\rho^{\pm}$ and $K^{\pm}=D^{\pm}$ for species). In the context of this given article, the capacities $C^{\pm}$ are assumed constant to maintain the linearity of the partial differential equations (PDEs). However, it is acknowledged that theoretically, it is possible to handle variable capacities, which would introduce non-linearity into the equations. This simplification is employed to make the mathematical analysis and solution of the PDEs more manageable. The symbol $r^{\pm}$ represents any volumetric source or sink term.
@@ 2.2 Interfacial conditions
The balance equation Eq. (7) is not sufficient to close the problem Eq. (3). The flux balance (7) must be complemented by one additional interfacial relation. In the context of conjugate heat transfer, for example, the continuity of the dependent variable is typically imposed,
@@
where $\operatorname{He}$ represents Henry’s law constant, which tends to $\operatorname{He}\to 1$ when the solubilities match across the interface cussler_diffusion_2009; Sander2023. Also note that both Eqs. (8) and (9) include discontinuities (non-zero right-hand side $f$) to accommodate source terms pollack_kapitza_1969.
The proposed framework can also accommodate more complex boundary conditions, such as nonlinearities (rate laws encountered in heterogeneous kinetics) and mixed boundary conditions (e.g. reservoir models with effective transfer coefficients). As highlighted below, this is possible because the intrinsic diffusion equations on each domain are solved simultaneously, as opposed to a single transport equation for an effective fluid that occupies the union of both domains as in the one-fluid approach. This alleviates the modeling issue of effective properties in interfacial cells and enables the use of different models in each phase (convection, diffusion models…).
@@ 3 Discrete geometric representation
A wide variety of techniques exist to represent interfaces. For simple objects (cuboids, cylinders, spheres, etc.), constructive solid geometry (CSG) is a convenient option laidlaw_constructive_1986. More complex geometries can be approximated using simplicial meshes or splines (e.g. Bézier curves, NURBS) or described implicitly by a level-set function osher_fronts_1988, in which a signed-distance field is sampled on the background mesh and interpolated between nodes. The choice of interface representation is closely tied to the numerical strategy used to solve the surrounding fields. The Arbitrary Lagrangian–Eulerian (ALE) formulation, for instance, is well suited when the interface is discretized with segments (in $d=2$) or triangles (in $d=3$) but it requires moving unstructured meshes. On fixed structured meshes (predominantly Cartesian), the flows on either side of the interface can instead be computed using the cut-cell method, a finite-volume approach typically classified as an embedded boundary method. This technique has been applied to the solution of the incompressible (single-phase) Navier-Stokes equations in complex geometries, first by Veldman & Verstappen verstappen_symmetry-preserving_2004 and later by Botella cheny_ls-stag_2010. In both cases, control volumes are explicitly constructed by determining the intersection of the interface with the Cartesian mesh. Once this construction is performed, differentiation operators are derived using the finite-volume methodology.
@@
For this reason, the proposed method does not rely on an explicit construction of trimmed polyhedra. Instead, it relies on Green-Gauss theorem and quadrature rules to derive the discrete operators. These quadrature rules require only a reduced set of geometric moments, described in Sec. 3.2. The computation of these moments thus constitutes the main bottleneck of the proposed method. It can be carried out with a handful of computational-geometry techniques and libraries. As shown below, this algorithmic choice is facilitated by the fact that all required moments reduce to intersections of hyperrectangles with the light (or dark) domain. For this purpose, efficient algorithms exist for this task. For implicitly defined geometries, Saye’s high-order quadrature methods provide accurate surface/volume integration directly from a level-set description saye_high-order_2015. The VOFI library also offers high-precision volume-of-fluid integration and related geometric quantities for implicitly defined interfaces chierici_optimized_2022. In two dimensions, general-purpose geometry engines such as GEOS/LibGEOS provide robust polygon operations (intersection, area, centroid)  santilli_geos_2023, which can be useful for handling complex planar geometries. These are also suitable, though less efficient owing to their generality. Altogether, such tools substantially simplify the implementation of the proposed method.
@@ 3.1 Geometrical foundations
Let’s now introduce all of geometric quantities required to assemble the discrete operators in Sec. 4. For the sake of brevity, this section focuses on the two-dimensional case ($d=2$). Three-dimensional extension can readily be performed.
@@
$\Omega_{i,j}=\Delta^{1}_{i}\times\Delta^{2}_{j},\quad\left(i,j\right)\in\left[1\mathinner{\ldotp\ldotp}N\right]\times\left[1\mathinner{\ldotp\ldotp}M\right]$
@@
$\Omega^{\pm}_{i,j}=\Omega_{i,j}\cap\Omega^{\pm}$ (10)
@@
$\Gamma_{i,j}=\Omega_{i,j}\cap\Gamma$
@@
where all the subsets on the right-hand side are disjoint represented in Fig 2(b).
@@
The description is completed by defining the corresponding faces in $\mathbb{R}$, that is defining both,
@@
The qualifiers employed earlier for the cells (pure and mixed) also apply to these faces. Finally, the subsets thus defined realize a partition of the boundary (vertices excluded) of each phase
@@ 3.2 Reduced geometric description
3.2.1 Geometric moments of the first kind
@@
The centroids along $x$ are defined as,
@@
The definitions (15) and (16) remain valid for empty cells, for which either $V_{i,j}^{+}$ or $V_{i,j}^{-}$ vanishes. In such instances, this enables a consistent definitions of the geometric moments of the second kind, defined in the next section.
@@
We also introduce the following face-centered quantity,
@@
Figure 5: Cut cell $\Omega_{i,j}$ with the domains $\Omega_{i,j}^{+}$ (dark) and $\Omega_{i,j}^{-}$ (light), with face‐areas $A^{1\pm}_{i\pm\frac{1}{2},j}$, $A^{2\pm}_{i,j\pm\frac{1}{2}}$. In relation to Fig. 3, $\Sigma$ denotes the topology of the embedded interface within the cell, whereas AA represents the face areas used in the flux discretization.
@@
3.2.2 Geometric moments of the second kind
The geometric moments defined below are referred to as “of the second kind” simply because their definition depends on the previously introduced “geometric moments of the first kind”.
We define the following cell-centered averages,
@@ 3.2.2 Geometric moments of the second kind
A summary table, located in the appendix 10, summarizes, for two spatial dimensions, the primary and secondary moments used in the discrete formulation.
@@ 4 Two-phase cut-cell method for static domains
where $t_{0}$ is arbitrarily set to 0.
@@
These interfacial unknowns are associated with the two interface conditions Eqs. (7) and (8).
@@ 4.1 Semi-discrete bulk equations
As for the primary variables, we represent the Cartesian components of the flux $\mathbf{q}^{\pm}$ by face averages. Let $q^{1\pm}:=\mathbf{q}^{\pm}\cdot\mathbf{e}^{1}$ and $q^{2\pm}:=\mathbf{q}^{\pm}\cdot\mathbf{e}^{2}$, where $\mathbf{e}^{m}$ denotes the $m$-th Cartesian basis vector. On vertical faces, we define
@@ 4.1.1 Volume-integrated divergence operator
Using the partition (12), we split the boundary integral into its mesh-face and interface contributions:
@@
$\left\{\begin{aligned} V^{\pm}_{i,j}\,\operatorname{div}^{1\omega}_{i,j}(A,Q)&:=A_{i+\nicefrac{{1}}{{2}},j}\,Q_{i+\nicefrac{{1}}{{2}},j}(t)-A_{i-\nicefrac{{1}}{{2}},j}\,Q_{i-\nicefrac{{1}}{{2}},j}(t),\\ V^{\pm}_{i,j}\,\operatorname{div}^{2\omega}_{i,j}(A,Q)&:=A_{i,j+\nicefrac{{1}}{{2}}}\,Q_{i,j+\nicefrac{{1}}{{2}}}(t)-A_{i,j-\nicefrac{{1}}{{2}}}\,Q_{i,j-\nicefrac{{1}}{{2}}}(t).\end{aligned}\right$. (28)
@@
The last term in (26), representing interfacial exchange, must be approximated using the available reduced geometric information. The discrete formula is designed to satisfy the following requirements: (i) minimal stencil, (ii) exact cancellation when the face fluxes are constant, (iii) at least first-order accuracy for mesh-aligned planar interfaces and (iv) recovery of the classical divergence away from interfaces.
@@
$\left\{\begin{aligned} V^{\pm}_{i,j}\,\operatorname{div}^{1\gamma}_{i,j}(A,B,Q)&:=\bigl(B_{i,j}-A_{i+\nicefrac{{1}}{{2}},j}\bigr)\,Q_{i+\nicefrac{{1}}{{2}},j}(t)+\bigl(A_{i-\nicefrac{{1}}{{2}},j}-B_{i,j}\bigr)\,Q_{i-\nicefrac{{1}}{{2}},j}(t),\\ V^{\pm}_{i,j}\,\operatorname{div}^{2\gamma}_{i,j}(A,B,Q)&:=\bigl(B_{i,j}-A_{i,j+\nicefrac{{1}}{{2}}}\bigr)\,Q_{i,j+\nicefrac{{1}}{{2}}}(t)+\bigl(A_{i,j-\nicefrac{{1}}{{2}}}-B_{i,j}\bigr)\,Q_{i,j-\nicefrac{{1}}{{2}}}(t).\end{aligned}\right$. (30)
Summing the bulk ($\omega$) and interfacial ($\gamma$) contributions (27) and (29), the terms proportional to $A^{m\pm}$ cancel and the divergence can be expressed using $B^{m\pm}$ only:
$\left\{\begin{aligned} V^{\pm}_{i,j}\,\operatorname{div}^{1}_{i,j}(B,Q)&:=B_{i,j}\,\bigl(Q_{i+\nicefrac{{1}}{{2}},j}(t)-Q_{i-\nicefrac{{1}}{{2}},j}(t)\bigr),\\ V^{\pm}_{i,j}\,\operatorname{div}^{2}_{i,j}(B,Q)&:=B_{i,j}\,\bigl(Q_{i,j+\nicefrac{{1}}{{2}}}(t)-Q_{i,j-\nicefrac{{1}}{{2}}}(t)\bigr).\end{aligned}\right$.

### v2

@@ Abstract
We present a Cartesian cut-cell finite-volume method for sharp-interface two-phase diffusion problems in static geometries. The formulation follows a two-fluid approach: independent diffusion equations are discretized in each phase on a fixed Cartesian grid, while the phases are coupled through embedded interface conditions enforcing continuity of diffusive flux and a general jump law. Cut cells are treated by integrating the governing equations over phase-restricted control volumes and surfaces, yielding discrete divergence and gradient operators that are locally conservative within each phase. Interface coupling is achieved by introducing a small set of interfacial unknowns per cut cell on the embedded boundary; the resulting algebraic system involves only bulk and interfacial averages. A key feature of the method is the use of a reduced set of geometric information based solely on low-order moments (trimmed volumes, apertures and interface measures/centroids), allowing robust implementation without constructing explicitly cut-cell polytopes. The method supports steady (Poisson) and unsteady (diffusion) regimes and incorporates Dirichlet, Neumann, Robin boundary conditions and general jumps. We validate the scheme on one-, two- and three-dimensional single-phase and two-phase benchmarks, including curved embedded boundaries, Robin conditions and strong property/jump contrasts. The results demonstrate a superlinear convergence behavior, sharp enforcement of interfacial laws and excellent conservation properties. Extensions to moving interfaces and Stefan-type free-boundary problems are natural perspectives of this framework.
@@ 1 Introduction
Diffusive transport across material interfaces is central to a wide range of heat- and mass-transfer phenomena, including conjugate heat transfer between solids and fluids, interphase species transfer, dissolution/precipitation in porous media and phase-change processes. At the continuum level, these problems are naturally formulated as two-phase diffusion models in which a scalar field (e.g. temperature or concentration) satisfies diffusion equations with phase-dependent coefficients and is coupled across a sharp interface through jump relations and flux balance (bird_transport_2007; cussler_diffusion_2009; carslaw_conduction_1980; Ishii2011). Depending on the application, the interface may enforce continuity of the scalar (conjugate heat transfer), a weighted jump such as Henry’s law (partitioning in mass transfer) or a finite interfacial resistance such as Kapitza resistance (Sander2023; pollack_kapitza_1969). These interfacial relations must be satisfied accurately to predict transfer rates, which are often limited by diffusive fluxes localized near the interface.
From a numerical standpoint, sharp-interface diffusion on complex geometries is challenging whenever the computational mesh does not conform to the interface. Body-fitted (ALE-type) approaches can represent the interface explicitly but require mesh generation and mesh motion, with the attendant complexity and robustness issues in three dimensions (hirt_arbitrary_1974). On fixed Cartesian grids, a broad class of non-conforming strategies has therefore been developed. Immersed-boundary and ghost/immersed-interface-type methods modify stencils near the interface or introduce forcing terms so that boundary and interface conditions are satisfied approximately, often with excellent accuracy for smooth solutions but with less direct control over strict local conservation (peskin_flow_1972; gibou_second-order-accurate_2002; gabbard_high-order_2024; verzicco_immersed_2023). Interface-capturing methods (VOF or level-set approaches) (hirt_volume_1981; osher_fronts_1988; popinet_accurate_2009) coupled with a one-fluid formulation offer flexibility for complex topological changes; however, enforcing two-phase diffusion jump conditions sharply on a fixed grid typically requires additional reconstruction and coupling machinery (and, in practice, often relies on one-fluid effective-property closures in interfacial cells). Hybrid combinations have recently been explored for solidification/melt problems (limare_hybrid_2023) and XFEM/level-set formulations have also been used for coupled Stefan/mass-transport settings (li_numerical_2019), highlighting continued interest in sharp and accurate interfacial transfer.
@@
  • a validation on one-, two- and three-dimensional benchmarks, mono- and two-phase, demonstrating a super-linear accuracy behavior together with conservation properties at the discrete level.
The paper is organized as follows. Sec. 2 introduces the two-phase continuum model and the interfacial jump conditions considered. Sec. 3 presents the embedded-boundary geometric setting and the reduced moment description used throughout. Sec. 4 derives the cut-cell finite-volume operators and the coupled discrete system. Numerical results are then reported in Sec. 5 to assess accuracy, robustness and conservation and Sec. 6 concludes and outlines extensions.
@@ 2 Continuum modeling
We introduce a generic scalar field $\phi$ (e.g. temperature $T$ or mass fraction $x$). The following describes the bulk transport model and the interfacial matching conditions used to close the two-phase diffusion problem.
@@ 2.1 Bulk transport equations
where $K^{\pm}$ is the diffusive mobility. The capacity $C^{\pm}$ collects the storage term (e.g., $C^{\pm}=\rho^{\pm}c_{p}^{\pm}$ and $K^{\pm}=k^{\pm}$ for heat; $C^{\pm}=\rho^{\pm}$ and $K^{\pm}=D^{\pm}$ for species). In the context of this given article, the capacities $C^{\pm}$ are assumed constant to maintain the linearity of the partial differential equations (PDEs). Variables capacities can also be treated but they introduce non-linearities which go beyond the scope or the current contribution. The symbol $r^{\pm}$ represents any volumetric source or sink term.
@@ 2.2 Interfacial conditions
The balance equation Eq. (7) is not sufficient to close the problem Eq. (3). The flux balance (7) must be complemented by one additional interfacial relation. In the context of conjugate heat transfer, for example, the jump of the dependent variable is typically imposed,
@@
where $\operatorname{He}$ represents Henry’s law constant, which tends to $\operatorname{He}\to 1$ when the solubilities match across the interface cussler_diffusion_2009; Sander2023. Also note that both Eqs. (8)-(9) include discontinuities (non-zero right-hand side $f$) to accommodate source terms pollack_kapitza_1969.
The proposed framework can also accommodate more complex boundary conditions, such as nonlinearities (rate laws encountered in heterogeneous kinetics) and mixed boundary conditions (e.g. reservoir models with effective transfer coefficients). These extensions are not discussed here for brevity. As highlighted below, this is possible because the intrinsic diffusion equations on each domain are solved simultaneously, as opposed to a single transport equation for an effective fluid that occupies the union of both domains as in the one-fluid approach. This alleviates the modeling issue of effective properties in interfacial cells and enables the use of different models in each phase (convection, diffusion models…).
@@ 3 Discrete geometric representation
A wide variety of techniques exist to represent interfaces. For simple objects (cuboids, cylinders, spheres, etc.), constructive solid geometry (CSG) is a convenient option laidlaw_constructive_1986. More complex geometries can be approximated using simplicial meshes or splines (e.g. Bézier curves, NURBS) or described implicitly by a level-set function osher_fronts_1988, in which a signed-distance field is either specified analytically or sampled on the background mesh and interpolated between nodes. The choice of interface representation is closely tied to the numerical strategy used to solve the surrounding fields. The Arbitrary Lagrangian–Eulerian (ALE) formulation, for instance, is well suited when the interface is discretized with segments (in $d=2$) or triangles (in $d=3$) but it requires moving unstructured meshes. On fixed structured meshes (predominantly Cartesian), the flows on either side of the interface can instead be computed using the cut-cell method, a finite-volume approach typically classified as an embedded boundary method. This technique has been applied to the solution of the incompressible (single-phase) Navier-Stokes equations in complex geometries on staggered meshes, first by Veldman & Verstappen verstappen_symmetry-preserving_2004 and later by Botella cheny_ls-stag_2010. In both cases, control volumes are explicitly constructed by determining the intersection of the interface with the Cartesian mesh. Once this construction is performed, differentiation operators are derived using the finite-volume methodology.
@@
For this reason, the proposed method does not rely on an explicit construction of trimmed polyhedra. Instead, it relies on Green-Gauss theorem and quadrature rules to derive the discrete operators. These quadrature rules require only a reduced set of geometric moments, described in Sec. 3.2. The computation of these moments thus constitutes the main bottleneck of the proposed method. It can be carried out with a handful of computational-geometry techniques and libraries. As shown below, this algorithmic choice is facilitated by the fact that all required moments reduce to intersections of hyperrectangles with the light (or dark) domain. For this purpose, efficient algorithms exist. For implicitly defined geometries, Saye’s high-order quadrature methods provide accurate surface/volume integration directly from a level-set description saye_high-order_2015. The VOFI library also offers high-precision volume-of-fluid integration and related geometric quantities for implicitly defined interfaces chierici_optimized_2022. In two dimensions, general-purpose geometry engines such as GEOS/LibGEOS provide robust polygon operations (intersection, area, centroid)  santilli_geos_2023, which can be useful for handling complex planar geometries. Altogether, such tools substantially simplify the implementation of the proposed method.
@@ 3.1 Geometrical foundations
We now introduce all of geometric quantities required to assemble the discrete operators in Sec. 4. For the sake of brevity, this section focuses on the two-dimensional case ($d=2$). Three-dimensional extension can readily be performed.
@@
$\Omega_{i,j}:=\Delta^{1}_{i}\times\Delta^{2}_{j},\quad\left(i,j\right)\in\left[1\mathinner{\ldotp\ldotp}N\right]\times\left[1\mathinner{\ldotp\ldotp}M\right]$
@@
$\Omega^{\pm}_{i,j}:=\Omega_{i,j}\cap\Omega^{\pm}$ (10)
@@
$\Gamma_{i,j}:=\Omega_{i,j}\cap\Gamma$
@@
where all the subsets on the right-hand side are disjoint (See Fig 2(b)).
@@
The description is completed by defining the corresponding faces in $\mathbb{R}^{2}$, that is defining both,
@@
The qualifiers employed earlier for the cells (pure and mixed) also apply to these faces. Finally, the subsets thus defined realize a partition of the boundary (vertices excluded) of each phase (See Fig 3),
@@ 3.2 Reduced geometric description
3.2.1 Primary geometric moments
@@
The centroids along $x$ are defined as (See Fig 4),
@@
The definitions Eq. (15) and Eq. (16) remain valid for empty cells, for which either $V_{i,j}^{+}$ or $V_{i,j}^{-}$ vanishes. In such instances, this enables a consistent definitions of the secondary geometric moments, defined in the next section.
@@
We also introduce the following face-centered quantity (See Fig 5),
@@
Figure 5: Cut cell $\Omega_{i,j}$ with the domains $\Omega_{i,j}^{+}$ (dark) and $\Omega_{i,j}^{-}$ (light), with face‐areas $A^{1\pm}_{i\pm\frac{1}{2},j}$, $A^{2\pm}_{i,j\pm\frac{1}{2}}$.
@@
3.2.2 Secondary geometric moments
The geometric moments defined below are referred to as “secondary” simply because their definition depends on the previously introduced “primary geometric moment”.
We define the following cell-centered averages (See Fig 6),
@@ 3.2.2 Secondary geometric moments
A summary table, located in the appendix 10, summarizes, for two spatial dimensions, the geometric moments used in the discrete formulation.
@@ 4 Two-phase cut-cell method for static domains
where $t_{0}$ is set to 0, consistently with the continuum model (See Sec. 2)
@@
These interfacial unknowns are associated with the two interface conditions (Eqs. (7)-(8) or Eqs. (7)-(9) depending on the the problem at hand).
@@ 4.1 Semi-discrete bulk equations
As for the primary variables, we represent the Cartesian components of the flux $\mathbf{q}^{\pm}$ by face averages. Let $q^{1\pm}:=\mathbf{q}^{\pm}\cdot\mathbf{e}^{1}$ and $q^{2\pm}:=\mathbf{q}^{\pm}\cdot\mathbf{e}^{2}$, where $\mathbf{e}^{\alpha}$ denotes the $\alpha$-th Cartesian basis vector. On vertical faces, we define
@@ 4.1.1 Volume-integrated divergence operator
Using the partition Eq. (12), we split the boundary integral into its mesh-face and interface contributions:
@@
$\left\{\begin{aligned} V^{\pm}_{i,j}\,\operatorname{div}^{1\omega}_{i,j}(A,Q)&:=A_{i+\nicefrac{{1}}{{2}},j}\,Q_{i+\nicefrac{{1}}{{2}},j}-A_{i-\nicefrac{{1}}{{2}},j}\,Q_{i-\nicefrac{{1}}{{2}},j},\\ V^{\pm}_{i,j}\,\operatorname{div}^{2\omega}_{i,j}(A,Q)&:=A_{i,j+\nicefrac{{1}}{{2}}}\,Q_{i,j+\nicefrac{{1}}{{2}}}-A_{i,j-\nicefrac{{1}}{{2}}}\,Q_{i,j-\nicefrac{{1}}{{2}}}.\end{aligned}\right$. (28)
@@
The last term in Eq. (26), representing interfacial exchange, must be approximated using the available reduced geometric information. The discrete formula is designed to satisfy the following requirements: (i) minimal stencil, (ii) exact cancellation when the face fluxes are constant, (iii) at least first-order accuracy for mesh-aligned planar interfaces and (iv) recovery of the classical divergence away from interfaces.
@@
$\left\{\begin{aligned} V^{\pm}_{i,j}\,\operatorname{div}^{1\gamma}_{i,j}(A,B,Q)&:=\bigl(B_{i,j}-A_{i+\nicefrac{{1}}{{2}},j}\bigr)\,Q_{i+\nicefrac{{1}}{{2}},j}+\bigl(A_{i-\nicefrac{{1}}{{2}},j}-B_{i,j}\bigr)\,Q_{i-\nicefrac{{1}}{{2}},j},\\ V^{\pm}_{i,j}\,\operatorname{div}^{2\gamma}_{i,j}(A,B,Q)&:=\bigl(B_{i,j}-A_{i,j+\nicefrac{{1}}{{2}}}\bigr)\,Q_{i,j+\nicefrac{{1}}{{2}}}+\bigl(A_{i,j-\nicefrac{{1}}{{2}}}-B_{i,j}\bigr)\,Q_{i,j-\nicefrac{{1}}{{2}}}.\end{aligned}\right$. (30)
Summing the bulk ($\omega$) and interfacial ($\gamma$) contributions Eq. (27) and Eq. (29), the terms proportional to $A^{\alpha\pm}$ cancel and the divergence can be expressed using $B^{\alpha\pm}$ only:
$\left\{\begin{aligned} V^{\pm}_{i,j}\,\operatorname{div}^{1}_{i,j}(B,Q)&:=B_{i,j}\,\bigl(Q_{i+\nicefrac{{1}}{{2}},j}-Q_{i-\nicefrac{{1}}{{2}},j}\bigr),\\ V^{\pm}_{i,j}\,\operatorname{div}^{2}_{i,j}(B,Q)&:=B_{i,j}\,\bigl(Q_{i,j+\nicefrac{{1}}{{2}}}-Q_{i,j-\nicefrac{{1}}{{2}}}\bigr).\end{aligned}\right$.